# The early history of symmetric teleparallel gravity: An overlooked period

### v1

@@ Abstract
It is noteworthy that symmetric teleparallel gravity has attracted considerable attention in recent years. A survey of the literature indicates that this surge of interest became particularly prominent around 2017 and 2018. However, together with my students and collaborators, we published a series of systematic and pioneering papers on this subject between 2004 and 2013. In this letter, we will summarize our overlooked contributions to symmetric teleparallel gravity produced during that period. For the sake of completeness and coherence, we will also briefly review our work on this topic carried out after 2018. In the final paragraph, we will write our personal perspectives on the future of symmetric teleparallel geometry.
@@
After this period, we suspended our research on STPG, since our earlier works received almost no attention within the alternative gravity community. Subsequently, beginning around 2017–2018, a series of influential papers on symmetric teleparallel gravity and later on $f(Q)$ gravity were produced by Tomi Koivisto, Alexey Golovnev, Jose Beltran Jimenez, Lavinia Heisenberg, and their collaborators. Following these developments, the number of studies in the field of STPG increased rapidly [13, 19, 22, 23] and the references therein, as well as the works citing them.
@@
Up to this point, we have summarized our contributions to the field of symmetric teleparallel gravity. More recently, we have been striving to apply the geometrical insight and computational techniques developed in this context to the theory of defects observed in crystalline structures. In the mainstream differential–geometric description of crystal defects, curvature is typically associated with disclinations, torsion with dislocations, and point defects with the trace of non-metricity. In contrast, we formulate all types of defects entirely within the framework of general teleparallel geometry [31, 32]. Using the model developed in this approach, we continue to work on projects addressing certain concrete problems studied in mechanical engineering, seeking symmetric teleparallel geometric solutions to them. We anticipate that symmetric teleparallel geometry will find increasing applications in physics, engineering, health sciences, and finance in the future, and we are actively pursuing research along these directions.
@@
Acknowledgements: This work has been supported by TÜBİTAK (The Scientific and Technological Research Council of Türkiye) via the project 124F325.

### v2

@@ Abstract
It is noteworthy that symmetric teleparallel gravity has attracted considerable attention in recent years. A survey of the literature indicates that this surge of interest became particularly prominent around 2017 and 2018. However, together with my students and collaborators, we published a series of systematic and pioneering papers on this subject between 2004 and 2013. This article aims to clarify the early geometric development of symmetric teleparallel gravity preceding the recent surge of interest in $f(Q)$ gravity. For the sake of completeness and coherence, we will also briefly review our work on this topic carried out after 2018. In the final paragraph, we will write our personal perspectives on the future of symmetric teleparallel geometry.
@@
The rapid development of symmetric teleparallel gravity after 2017–2018 has created a substantial body of literature whose conceptual roots extend to earlier geometric formulations. A systematic reconstruction of this early stage is therefore useful for establishing continuity in the development of the subject and for clarifying the origin of several structures that reappear in modern formulations.
@@
After this period, we suspended our research on STPG, since our earlier works received almost no attention within the alternative gravity community. Subsequently, beginning around 2017–2018, a series of influential papers on symmetric teleparallel gravity and later on $f(Q)$ gravity were produced by Tomi Koivisto, Alexey Golovnev, Jose Beltran Jimenez, Lavinia Heisenberg, and their collaborators. The recent resurgence of interest in symmetric teleparallel formulations, particularly through coincident general relativity and $f(Q)$ gravity, motivates a reexamination of earlier geometric approaches. Following these developments, the number of studies in the field of STPG increased rapidly [13, 19, 21, 22] and the references therein, as well as the works citing them.
@@
Up to this point, we have summarized our contributions to the field of symmetric teleparallel gravity. More recently, we have been striving to apply the geometrical insight and computational techniques developed in this context to the theory of defects observed in crystalline structures. In the mainstream differential–geometric description of crystal defects, curvature is typically associated with disclinations, torsion with dislocations, and point defects with the trace of non-metricity. In contrast, we formulate all types of defects entirely within the framework of general teleparallel geometry [30, 31]. Using the model developed in this approach, we continue to work on projects addressing certain concrete problems studied in mechanical engineering, seeking symmetric teleparallel geometric solutions to them. We anticipate that symmetric teleparallel geometry will find increasing applications in physics, engineering, health sciences, and finance in the future, and we are actively pursuing research in these directions.
The purpose of this article is, therefore, not merely historical documentation but the clarification of the conceptual lineage of symmetric teleparallel geometry and its relation to ongoing research programs.
@@
Acknowledgements: This work has been supported by TÜBİTAK (The Scientific and Technological Research Council of Türkiye) with the grant id 124F325.